Find the total surface area in square metres of the outside of an open pipe of radius 85 cm and length 4.5m, correct to two decimal places
Answer by rothauserc(4718)

surface area(SA) of open pipe = 2 * pi * r * l where pi = 22/7, r is the pipe's radius, pipe's length
:
SA = 2 * (22/7) * (0.85) * 4.5
:
***************************************
SA = 24.0429 approx 24.04 square metres
***************************************
: